OpenStudy (jiteshmeghwal9):

\[\frac{(\log_{}{x}-\log_{}{y} )(\log_{}{x^2}+\log_{}{y^2} )}{(\log_{}{x^2}-\log_{}{y^2} )(\log_{}{x+y} )}\] is equal to????

OpenStudy (foolaroundmath):

\[\log{x} + \log{y} = \log{xy}\]\[\log{x}-\log{y} = \log\frac{x}{y}\] Is the denominator \(\log{x}+\log{y} \)? instead of \(\log{x}+y\)

OpenStudy (anonymous):

\[\frac{\log(\frac{x}{y}) \times \log(x^2y^2)}{\log(\frac{x^2}{y^2}) \times \log(xy)} \implies \frac{\log(\frac{x}{y}) \times 2\log(xy)}{2\log(\frac{x}{y}) \times \log(xy)} \implies 1\]

OpenStudy (anonymous):

\[\large \log{x} + \log{y} = \log{xy}\] \[\large \log{x}-\log{y} = \log\frac{x}{y}\] I have used this formulas in first step.. \[\large loga^b = b \times \log(a)\] In second step I have used this...

OpenStudy (jiteshmeghwal9):

i used all this formulas but i didn't gt the answer @waterineyes :(

OpenStudy (anonymous):

I have solved it answer is 1..

OpenStudy (jiteshmeghwal9):

\[\LARGE{Ans:1}\]

OpenStudy (jiteshmeghwal9):

so plz give me full solution

OpenStudy (anonymous):

See above i have solved it..

OpenStudy (jiteshmeghwal9):

oh! ya i gt it. Thanx a lot:)

OpenStudy (anonymous):

Welcome dear..